# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

### v3

@@ 4.3 Ablation study
A challenging aspect of our problem setting is that we do not know a priori the best way(s) to divide the $T$ objectives into $K$ subsets such that good coverage can be obtained. In this section, we seek to answer the question: what is the efficiency lost by MOCOBO due to not knowing an efficient partitioning of the objectives in advance?. To construct a proxy “efficient” partition to measure this, we first run MOCOBO to completion on a task, and then consider optimization as if we had known the partitioning of objectives found at the end in advance.
@@
With such an “oracle” partitioning in hand ahead of time, we can efficiently optimize by running $K$ independent optimization runs: one to find a single solution for each subset in the fixed, given partition. With a fixed partition, the coverage score in (1) reduces to optimizing the sum of objectives in each relevant subset independently. In Figure˜3 right, we compare this strategy directly to MOCOBO on three optimization tasks. We plot the best coverage score obtained by $K$ independent runs of TuRBO/LOL-BO on the fixed oracle $K$-partitioning of the objectives by combining their solutions into a single set. MOCOBO achieves the same average best coverage scores with minimal loss in optimization efficiency despite having to discover an efficient partitioning during optimization. See Section˜A.2.2 for an additional ablation study of MOCOBO when all objectives are highly conflicting.
5 Discussion
@@
By bridging the gap between traditional multi-objective Bayesian optimization (BO) and practical coverage requirements, MOCOBO offers an effective approach to tackle critical problems in drug design and beyond. This framework extends the reach of Bayesian optimization into new domains, providing a robust solution to the challenges posed by extreme trade-offs and the need for specialized, collaborative solutions. See Section˜A.4 for additional discussion and limitations of MOCOBO.

### v4

@@ 4.3 Ablation study
A challenging aspect of our problem setting is that we do not know a priori the best way(s) to divide the $T$ objectives into $K$ subsets such that good coverage can be obtained. In this section, we seek to answer the question: what is the efficiency lost by MOCOBO due to not knowing an efficient partitioning of the objectives in advance? To construct a proxy “efficient” partition to measure this, we first run MOCOBO to completion on a task, and then consider optimization as if we had known the partitioning of objectives found at the end in advance.
@@
With such an “oracle” partitioning in hand ahead of time, we can efficiently optimize by running $K$ independent optimization runs: one to find a single solution for each subset in the fixed, given partition. With a fixed partition, the coverage score in (2) reduces to optimizing the sum of objectives in each relevant subset independently. In Figure˜4, we compare this strategy directly to MOCOBO on three optimization tasks. We plot the best coverage score obtained by $K$ independent runs of TuRBO/LOL-BO on the fixed oracle $K$-partitioning of the objectives by combining their solutions into a single set. MOCOBO achieves the same average best coverage scores with minimal loss in optimization efficiency despite having to discover an efficient partitioning during optimization. See Section˜B.2 for additional ablation studies.
5 Related Works
The coverage optimization problem setting considered in this paper has not been previously explored in the Bayesian optimization literature. However, similar problem settings [17, 18, 19] have been studied in the context of gradient-based multi-objective optimization. For instance, Ding et al. [17] demonstrated that the coverage problem generalizes various clustering methods, including k-means clustering, and proposed an algorithm that extends k-means++ [58] and Lloyd’s algorithm [59] to efficiently solve the coverage problem using gradient descent. Li et al. [18] introduced Many-objective Multi-solution Transport (MosT), a novel framework that scales multi-objective gradient-based optimization to a large number of objectives. MosT uses an optimal transport to establish a balanced assignment between objectives and solutions, allowing each solution to focus on a specific subset of objectives and ensuring that the collective set of solutions covers all objectives. Lin et al. [19] introduced a novel Tchebycheff set (TCH-Set) scalarization approach for covering multiple conflicting objectives using a small set of collaborative solutions. They further propose a smooth Tchebycheff set (STCH-Set) scalarization method to handle non-smoothness in TCH-Set scalarization. The reliance on gradients makes these previous approaches inapplicable to the black-box setting.
For the black-box setting, Liu et al. [20] was the first to consider the coverage problem. They propose an analogous formalization of the coverage problem in the context of multi-objective black-box optimization, along with a clustering-based swarm optimization algorithm (CluSO) designed to solve it. In Section˜4, we provide empirical results demonstrating that our Bayesian optimization approach (MOCOBO) consistently outperforms CluSO by a large margin across all tasks considered.
Lastly, we note that Malkomes et al. [29] proposed a method with the same name as the ECI we defined in Section˜3, which was later extended in [60]. The resemblance is only in the name, as their work addresses a fundamentally different notion of coverage. Mainly, they attempt to cover the input space by identifying a diverse set of feasible solutions that span a subspace defined by known threshold constraints on each objective. In contrast, we attempt to “cover” the different objectives.
6 Conclusions
@@
By bridging the gap between traditional multi-objective Bayesian optimization (BO) and practical coverage requirements, MOCOBO offers an effective approach to tackle critical problems in drug design and beyond. This framework extends the reach of Bayesian optimization into new domains, providing a robust solution to the challenges posed by extreme trade-offs and the need for specialized, collaborative solutions. See Appendix˜D for additional discussion and limitations of MOCOBO.